The postal service of St. Vincent, an island in the West Indies, obtains a significant portion of its revenues from sales of special souvenir sheets to stamp collectors. The souvenir sheets usually contain several high-value St. Vincent stamps depicting a common theme, such as the life of Princess Diana. The souvenir sheets are designed and printed for the postal service by Imperial Printing, a stamp agency service company in the United Kingdom. The souvenir sheets cost the postal service $1.25 each. St. Vincent has been selling these souvenir sheets for $8.00 each and ordinarily sells about 63,000 units. To test the market, the postal service recently priced a new souvenir sheet at $9.00 and sales dropped to 53,000 units.

2. Estimate the price elasticity of demand for the souvenir sheets. (Negative amount should be indicated by a minus sign. Do not round intermediate calculations. Round your answer to 4 decimal places.)


3. Estimate the profit-maximizing price for souvenir sheets. (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

2)

Given

P1=$8, Q1=63000

P2=$9, Q2=53000

I shall use mid point approach for calculating price elasticity of demand

Percent change in price=(9-8)/((9+8)/2)=11.764706%

Percent change in quantity demanded=(53000-63000)/((53000+63000)/2)=-17.241379%

Price elasticity of demand=Percent change in qunatity demanded/Percent change in price

Price elasticity of demand=-17.241379%/11.764706%=-1.465517 or say -1.4655

3)

Marginal cost =MC=$1.25

Profit maximizing price is give as

MC Pa 1+1

1.25 P=14145517 = $3.94

Profit maximizing price is $3.94